QUANTITATIVE ROTATIONAL SPECTROSCOPY FOR HITRAN

\underline{BRIAN J. DROUIN}, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109-8099.

Remote sensing of rotational transitions in the Earth’s atmosphere has become a popular method of retrieving geophysical temperatures, pressures and chemical composition profiles. Reliable retrieval of this information requires accurate spectral information. The HITRAN database is a repository for spectral information that is useful to the atmospheric remote sensing community. A majority of the rotational data in HITRAN are based on the JPL spectral line catalog and then augmented with lineshape parameters and references. This presentation outlines the current status of rotational data that are useful for atmospheric measurements.